# Weyl law for 1-cycles

### v1

@@ 3.3. Filling small families of cycles
where $V(\operatorname{Center}(E))=\operatorname{Center}(E)\cap C(3)_{0}$ is the set of vertices of $\operatorname{Center}(E)$. Hence it suffices to define $\tau_{0}$ on $V(\operatorname{Center}(E))$ for each face $E$ of $C$. Given a face $E$ of $C$, we choose $v_{E}\in V(E)$ (arbitrary) and we define $\tau_{0}(w)=\tau_{0}(v_{E})+\tau(\Xi(w),v_{E})$ where given $v,v^{\prime}\in V(E)$ we select $\tau(v,v^{\prime})\in\mathcal{I}_{k+1}(M)$ a chain such that $\partial\tau(v,v^{\prime})=F(v)-F(v^{\prime})$, $\operatorname{spt}(\tau(v,v^{\prime}))\subseteq U_{E}$ and $\operatorname{\mathbf{M}}(\tau(v,v^{\prime}))\leq 2\varepsilon$. Notice that $\operatorname{\mathbf{M}}(\tau_{0}(w))\leq 3\varepsilon$ for every $w\in C(3)_{0}$ and that in fact this definition is independent of the cell $C$ (we get the same $\tau_{0}$ if we regard $E$ as a face of another top-dimensional cell $C^{\prime}$) hence it extends $\tau_{0}$ to $X(3)_{0}$. But this refinement has some extra advantages. Given $x,y\in X(3)_{0}$ with $d(x,y)=1$ (here $d=d^{3}$ as in Definition 3.21), either there exists $v$ such that $x,y\in C_{v}$ and hence $\partial\tau_{0}(x)=\partial\tau_{0}(y)$ or $\tau_{0}(x)-\tau_{0}(y)$ is supported in $U_{E}$ because $x,y\in\operatorname{Center}(E)$ for some face $E$ of $C$. Thus to obtain $\delta$-localization we will perform refinements to the domain $\tilde{X}$ to replace “jumps” as in the first case from a chain $\tau_{0}(x)$ to another chain $\tau_{0}(y)$ with (possibly) very different support (but the same boundary) by “smaller jumps” involving consecutive chains with similar supports, taking advantage of the fact that $\tau(x)-\tau(y)$ is a $(k+1)$-cycle which we can chop using Lemma 3.23. We proceed by induction in the skeleta of $\tilde{X}$.
Inductive property 2. Given $1\leq j\leq p$, denote $q_{j}=q(k,j,p,M,\delta)$. There exists a discrete family $\tau_{j}:\tilde{X}(q_{j})_{0}\to\mathcal{I}_{k+1}(M)$ and for each $j$-cell $F$ of $\tilde{X}_{j}(q_{j})$ an $(K(N,k,j),K(N,k,j)\delta)$-admissible family $\{B_{i}^{F}\}_{i\in I_{F}}$ such that
@@
    If $F\sim F^{\prime}$ then $\{B_{i}\}_{i\in I_{F}}=\{B_{i}\}_{i\in I_{F^{\prime}}}$. Furthermore, let $\Delta^{j}_{FF^{\prime}}(x)=\tau_{j}(x)-\tau_{j}(T_{FF^{\prime}}(x))$ for $x\in F\cap\tilde{X}(q_{j})_{0}$ then $\Delta^{j}_{FF^{\prime}}(x)$ is supported in $\bigcup_{i\in I_{F}}B_{i}^{F}$.
Base case 2. Let $q_{1}=q_{1}(k,1,p,M,\delta)$ be the smallest odd number which is greater or equal than $L$. Define $\tau_{1}(x)=\tau_{0}(x)$ for $x\in\tilde{X}_{0}$. Our goal is to extend $\tau_{1}$ to $\tilde{X}_{1}(q_{1})_{0}$. To do that, we pick for each class $\mathcal{E}$ of $1$-cells under $\sim$ a representative $E\in\mathcal{E}$. We first define $\tau_{1}$ on $E$ and then we extend it to the other edges $E^{\prime}\in\mathcal{E}$. Let $C\in\operatorname{Faces}_{p}(X)$ and $v\in V(C)$ be such that $E\subseteq C_{v}$. Let $v_{E},w_{E}$ be the vertices of $E$. Observe that $\tau_{1}(w_{E})-\tau_{1}(v_{E})$ is a $(k+1)$-cycle because $\partial\tau_{1}(v_{E})=F(v)=\partial\tau_{1}(w_{E})$. Let $\sigma_{E}\in\mathcal{I}_{k+2}(M)$ be such that $\partial\sigma_{E}=\tau_{1}(w_{E})-\tau_{1}(v_{E})$ and $\operatorname{\mathbf{M}}(\sigma_{E})\leq(6\varepsilon)^{\frac{k+2}{k+1}}\leq 6\varepsilon$ (such $\sigma_{E}$ exists by the Isoperimetric Inequality). Using the Coarea Inequality, we can define for each $1\leq l\leq L$ a radius $r_{l}^{E}\in[r,2r]$ such that
$\operatorname{\mathbf{M}}(\sigma_{E}\llcorner\partial B(x_{l},r_{l}^{E}))\leq\frac{\operatorname{\mathbf{M}}(\sigma_{E})}{r}\leq 6\frac{\varepsilon}{r}.$
@@
where $l(x)=d(x,v_{E})$ (here $d$ denotes the metric $d^{q_{1}}$ on $\tilde{X}$ as in Definition 3.21). Then by construction, if $x,y\in E\cap\tilde{X}(q_{1})_{0}$ verify $d(x,y)=1$ then $\tau_{1}(x)-\tau_{1}(y)$ is supported in a certain ball $B(x_{l},2r)$ and
$\operatorname{\mathbf{M}}(\tau_{1}(x))\leq 3\varepsilon+6L\frac{\varepsilon}{r}\leq 12L\frac{\varepsilon}{r}.$
@@
so that we have $\partial\tau_{1}(x)=R^{q_{1}}F(x)$. For each $1$-cell $F$ of $\tilde{X}_{1}(q_{1})$ contained in $E$, let $l=d_{\infty}(F,\{v_{E}\})$ and let $\{B_{i}^{F}\}_{i\in I_{F}}$ be an $(2N+1,\delta_{1})$-admissible family which covers $B(x_{l+1},2\delta)\cup\bigcup_{i\in I_{E_{1}}}U_{i}^{E_{1}}\cup\bigcup_{i\in I_{E_{1}}}U_{i}^{E_{2}}$, which exists by Lemma 3.5 for $\delta_{1}=12\delta$. For $F^{\prime}=T_{EE^{\prime}}(F)$, define $\{B_{i}^{F^{\prime}}\}_{i\in I_{F^{\prime}}}=\{B_{i}^{F}\}_{i\in I_{F}}$. Then we have $\operatorname{\mathbf{M}}(\tau_{1}(x))\leq 18L\frac{\varepsilon}{\delta}$ (here we make use of our choice $r=\delta$) and if $x,y\in F^{\prime}$ then
@@
is supported in $\bigcup_{i\in I_{F^{\prime}}}B_{i}^{F^{\prime}}$. In case $|\mathcal{E}|=1$, we just define the admissible family to be $\{B(x_{l},2\delta)\}$. Thus properties (A1) to (A4) hold with $c(k,1)=18$, $n(k,1)=1$, and $K(N,k,1)=12$.
Inductive step 2. Assume $\tau_{j-1}$ has been defined in $\tilde{X}_{j-1}(q_{j-1})_{0}$ ($q_{j-1}=q(k,j-1,M,\delta)$) verifying properties (A1) to (A4). Define
$d_{j}$ $=q(k+1,j-1,M,\delta),$
@@
and for each $j$-cell $F\subseteq E$ of $X_{j}(q_{j})$ the family $\tau_{j}:V(F)\to\mathcal{Z}_{k}(M)$ is localized in a $(K_{2}(N,k,j),K_{2}(N,k,j)\delta)$ admissible family by Lemma 3.23 with $r=\delta$ which yields $K_{2}(N,k,j)=12(K_{1}(N,k,j)+1)$. If the equivalence class $\mathcal{E}$ of $E$ has more than one element, despite of the choice of $v_{E}$ there exists a unique cell $C_{E}$ of $X$ such that $v_{E}\in\operatorname{Center}(C_{E})$ and it holds that $T_{EE^{\prime}}(v_{E})\in\operatorname{Center}(C_{E})$ for every $E^{\prime}\in\mathcal{E}$. Let us enlarge the previously constructed admissible family to a $(K_{2}(N,k,j)+N,3(K_{2}(N,k,j)+1)\delta)$-localized family which also covers $\{U_{i}^{C_{E}}:i\in I_{C_{E}}\}$ (using Lemma 3.5). This will be the family $\{B_{i}\}_{i\in I_{F}}$ and thus we set $K(N,k,j)=3(K_{2}(N,k,j)+1)$. Notice that $\tau_{0}(v_{E})-\tau_{0}(T_{EE^{\prime}}(v_{E}))$ is supported in $\bigcup_{i\in I_{C_{E}}}U_{i}^{C_{E}}$ and therefore in $\{B_{i}\}_{i\in I_{F}}$ for every $E^{\prime}\sim E$.
Define $\tau_{j}(x)=R^{d_{j}}\tau_{j-1}(\Xi(x))$ for each $x\in X_{j-1}\cap X(q_{j})_{0}$, which extends the previous definition of $\tau_{j}$ along $\partial E$. For $x\in X_{j-1}\cap X(q_{j})_{0}\cap E$ set $\Delta_{EE^{\prime}}^{j}(x)=\tau_{j}(x)-\tau_{j}(T_{EE^{\prime}}(x))$. Let $E^{\prime}\in\mathcal{E}$, $E^{\prime}\neq E$. We extend $\Delta^{j}_{EE^{\prime}}(x)$ to $\operatorname{int}(E)\cap\tilde{X}(q_{j})_{0}$ by
@@
$\tau_{j}(x)=\tau_{j}(T_{E^{\prime}E}(x))-\Delta^{j}_{EE^{\prime}}(x)$
@@
$\operatorname{\mathbf{M}}(\tau_{j}(x))\leq\operatorname{\mathbf{M}}(\tau_{j}(T_{EE^{\prime}}(x)))+\operatorname{\mathbf{M}}(\Delta^{j}_{EE^{\prime}}(x))\leq c(k,j)(\frac{L}{\delta})^{n(k,j)}\varepsilon$
if we set $c(k,j)=9c(k+1,j-1)c(k,j-1)$ and $n(k,j)=n(k+1,j-1)+n(k,j-1)+1$. Additionally, if $x,y\in F^{\prime}\cap\tilde{X}(q_{j})_{0}$ then
$\tau_{j}(x)-\tau_{j}(y)=\tau_{j}(T_{E^{\prime}E}(x))-\tau_{j}(T_{E^{\prime}E}(y))+\Delta^{j}_{EE^{\prime}}(x)-\Delta^{j}_{EE^{\prime}}(y)$
which by construction of $\tau_{j}$ along $F$ is supported in $\bigcup_{i\in I_{F}}B_{i}^{F}$. In case $\mathcal{E}$ has only one element, the $(K_{2}(N,k,j),K_{2}(N,k,j)\delta)$-admissible families constructed first already have the desires properties. ∎
@@ 3.4. Approximating discrete families of cycles
There exist constants $\varepsilon_{0}(k,p,M)>0$, $\delta_{0}(k,p,M)>0$, $q(k,p,M,\delta)\in\mathbb{N}$ and $N(k,p)\in\mathbb{N}$ such that the following is true. Let $F:X_{0}\to\mathcal{Z}_{k}(M)$ be an $\varepsilon$-fine family, $\varepsilon\leq\varepsilon_{0}$, $10\leq k\leq n-$ 1, $\dim(X)=p$. Let $0<\delta<\delta_{0}$. Then there exists a discrete family $F^{\prime}:X(q)_{0}\to\mathcal{Z}_{k}(M)$ such that
@@
    $F^{\prime}$ is $c(k,p)(\frac{L}{\delta})^{n(k,p)}\varepsilon$-fine and $(N(k,p),\delta)$-localized.
We start by fixing for each pair of adjacent vertices $v,w\in X_{0}$ a $(k+1)$-chain $\tau_{vw}=\tau(v,w)$ such that $\operatorname{\mathbf{M}}(\tau_{vw})\leq\varepsilon$ and $\partial\tau_{vw}=F(v)-F(w)$. We fix points $p_{1},...,p_{L}$ in the manifold $M$ such that $\{B(p_{l},r):1\leq l\leq L\}$ covers $M$ ($r$ to be chosen in terms of $\delta$). They can be chosen so that $Lr^{n}\leq c(n)\operatorname{Vol}(M,g)$ for some universal constant $c(n)$ which depends only on the dimension of the manifold $M$. For each cell $C$ of $X$ we denote $V(C)=C\cap X_{0}$ the set of its vertices and we choose a preferred vertex $v_{C}\in V(C)$.
Following the strategy in [12], we will inductively define for each $1\leq j\leq p=\dim(X)$ a $\delta_{j}$-localized and $\varepsilon_{j}$-fine map $F_{j}:X_{j}(q_{j})_{0}\to\mathcal{Z}_{k}(M)$, where $q_{1}<...<q_{p}=q$. All of the previous maps will agree with $F$ at $X_{0}$. Given a $j$-dimensional cell $C$ of $X(q)$, the map $F_{j}|_{C\cap X(q_{j})_{0}}$ will be given by an interpolation of the values of $F(v)$ over the vertices $v$ of $C$ which does not increase the mass too much with respect to $\max\{\operatorname{\mathbf{M}}(F(v)):v\in C\cap X_{0}\}$. Let us explain how this interpolation works for $j=1$.
Let $C\in\operatorname{Faces}_{1}(X)$ and let $v,w$ be the vertices of $C$. We are going to choose a certain grid $\mathcal{D}_{C}=\{D_{1},...,D_{L}\}$ on $M$ consisting of $L$ domains with disjoint interiors and piecewise smooth boundaries. Roughly speaking, the idea to interpolate between $F(v)$ and $F(w)$ is to start from $F(v)$ and replace $F(v)\llcorner D_{l}$ by the chain with smallest mass among $F(w)\llcorner D_{l}$ and $F(v)\llcorner D_{l}$ for each $l=1,...,L$. This operation is done one domain at a time. Thus we start from $z_{0}=F(v)$ and inductively define $z_{l}=z_{l-1}$ if $\operatorname{\mathbf{M}}(F(v)\llcorner D_{l})\leq\operatorname{\mathbf{M}}(F(w)\llcorner D_{l})$ and
$z_{l}=z_{l-1}+\partial[\tau_{vw}\llcorner D_{l}]$
@@
otherwise. It can be shown that $\tilde{z}_{L}=z_{L}$ and hence this allows to define $F_{1}$ along $C(2l)_{0}$ by $z_{0},z_{1},...,z_{L},\tilde{z}_{L-1},...,\tilde{z}_{1},\tilde{z}_{0}$. This has the desired $\delta$-localization property as $z_{l}-z_{l-1}$ is supported in $D_{l}\subseteq B(x_{l},2r)$ for every $l$ (respectively $\tilde{z}_{l}-\tilde{z}_{l-1}$). We can formalize the previous in the following way. For each $x\in C\cap X(3L)_{0}$ and each $D\in\mathcal{D}_{C}$, we have a certain vertex $w^{D}_{C}(x)\in\{v,w\}$ such that
(2) $F_{1}(x)=F(v)+\sum_{D\in\mathcal{D}_{C}}\partial[\tau(v,w_{C}^{D}(x))\llcorner D]$
@@
for some family $I_{j}:X_{j}(q_{j})_{0}\to\mathcal{I}_{k}(M)$ with small mass in case we try to generalize (3). It will be convenient to work with (5) because when proceeding inductively from the $(j-1)$-skeleton to the $j$-skeleton, the preferred vertex $v_{E}$ of a $(j-1)$-face $E$ of $C$ will not always coincide with $v_{C}$ (for example when $v_{C}\notin E)$ and therefore we would need to do a certain adjustment to the $\tau(v_{E},w_{j}^{D}(x))$ in order to have (4) which is not necessary if we adopt (5). This problem already arises when $j=2$, as for $x\in E\subseteq\partial C$ with $v_{E}\neq v_{C}$ we would have

### v2

@@ 3.3. Filling discrete families of small cycles
where $V(\operatorname{Center}(E))=\operatorname{Center}(E)\cap C(3)_{0}$ is the set of vertices of $\operatorname{Center}(E)$. Hence it suffices to define $\tau_{0}$ on $V(\operatorname{Center}(E))$ for each face $E$ of $C$. Given a face $E$ of $C$, we choose $v_{E}\in V(E)$ (arbitrary) and we define $\tau_{0}(w)=\tau_{0}(v_{E})+\tau(\Xi(w),v_{E})$ where given $v,v^{\prime}\in V(E)$ we select $\tau(v,v^{\prime})\in\mathcal{I}_{k+1}(M)$ a chain such that $\partial\tau(v,v^{\prime})=F(v)-F(v^{\prime})$, $\operatorname{spt}(\tau(v,v^{\prime}))\subseteq U_{E}$ and $\operatorname{\mathbf{M}}(\tau(v,v^{\prime}))\leq 2\varepsilon$. Notice that $\operatorname{\mathbf{M}}(\tau_{0}(w))\leq 3\varepsilon$ for every $w\in C(3)_{0}$ and that in fact this definition is independent of the cell $C$ (we get the same $\tau_{0}$ if we regard $E$ as a face of another top-dimensional cell $C^{\prime}$) hence it extends $\tau_{0}$ to $X(3)_{0}$. But this refinement has some extra advantages. Given $x,y\in X(3)_{0}$ with $d(x,y)=1$ (here $d=d^{3}$ as in Definition 3.18), either there exists $v$ such that $x,y\in C_{v}$ and hence $\partial\tau_{0}(x)=\partial\tau_{0}(y)=F(v)$ or $\tau_{0}(x)-\tau_{0}(y)$ is supported in $U_{E}$ because $x,y\in\operatorname{Center}(E)$ for some face $E$ of $C$. Thus to obtain $\delta$-localization we will perform refinements to the domain $\tilde{X}$ to replace “jumps” as in the first case from a chain $\tau_{0}(x)$ to another chain $\tau_{0}(y)$ with (possibly) very different support but the same boundary by “smaller jumps” involving consecutive chains with similar supports, taking advantage of the fact that $\tau(x)-\tau(y)$ is a $(k+1)$-cycle which we can chop using Lemma 3.20. We proceed by induction in the skeleta of $\tilde{X}$.
Inductive property 2. Given $1\leq j\leq p$, denote $q_{j}=q(k,j,p,M,\delta)$. There exists a discrete family $\tau_{j}:\tilde{X}(q_{j})_{0}\to\mathcal{I}_{k+1}(M)$ and for each $j$-cell $F$ of $\tilde{X}_{j}(q_{j})$ a $(K(N,k,j),K(N,k,j)\delta)$-admissible family $\{B_{i}^{F}\}_{i\in I_{F}}$ such that
@@
    If $F\sim F^{\prime}$ then $\{B_{i}\}_{i\in I_{F}}=\{B_{i}\}_{i\in I_{F^{\prime}}}$. Furthermore, let $\Delta^{j}_{FF^{\prime}}(x)=\tau_{j}(x)-\tau_{j}(T_{FF^{\prime}}(x))$ for $x\in F\cap\tilde{X}(q_{j})_{0}$. Then $\Delta^{j}_{FF^{\prime}}(x)$ is supported in $\bigcup_{i\in I_{F}}B_{i}^{F}$.
Base case 2. Let $q_{1}=q_{1}(k,1,p,M,\delta)$ be the smallest odd number which is greater or equal than $L$. Define $\tau_{1}(x)=\tau_{0}(x)$ for $x\in\tilde{X}_{0}$. Our goal is to extend $\tau_{1}$ to $\tilde{X}_{1}(q_{1})_{0}$. To do that, we pick for each class $\mathcal{E}$ of $1$-cells under $\sim$ a representative $E\in\mathcal{E}$. We first define $\tau_{1}$ on $E$ and then we extend it to the other edges $E^{\prime}\in\mathcal{E}$. Let $C\in\operatorname{Faces}_{p}(X)$ and $v\in V(C)$ be such that $E\subseteq C_{v}$. Let $v_{E},w_{E}$ be the vertices of $E$. Observe that $\tau_{1}(w_{E})-\tau_{1}(v_{E})$ is a $(k+1)$-cycle because $\partial\tau_{1}(v_{E})=F(v)=\partial\tau_{1}(w_{E})$. Let $\sigma_{E}\in\mathcal{I}_{k+2}(M)$ be such that $\partial\sigma_{E}=\tau_{1}(w_{E})-\tau_{1}(v_{E})$ and $\operatorname{\mathbf{M}}(\sigma_{E})\leq(6\varepsilon)^{\frac{k+2}{k+1}}\leq 6\varepsilon$ (such $\sigma_{E}$ exists by the Isoperimetric Inequality). Using the Coarea Inequality, we can define for each $1\leq l\leq L$ a radius $r_{l}^{E}\in[\delta,2\delta]$ such that
$\operatorname{\mathbf{M}}(\sigma_{E}\llcorner\partial B(x_{l},r_{l}^{E}))\leq\frac{\operatorname{\mathbf{M}}(\sigma_{E})}{\delta}\leq 6\frac{\varepsilon}{\delta}.$
@@
where $l(x)=d(x,v_{E})$ (here $d$ denotes the metric $d^{q_{1}}$ on $\tilde{X}$ as in Definition 3.18). Then by construction, if $x,y\in E\cap\tilde{X}(q_{1})_{0}$ verify $d(x,y)=1$ then $\tau_{1}(x)-\tau_{1}(y)$ is supported in a certain ball $B(x_{l},2\delta)$ and
$\operatorname{\mathbf{M}}(\tau_{1}(x))\leq 3\varepsilon+6L\frac{\varepsilon}{\delta}\leq 9L\frac{\varepsilon}{\delta}.$
@@
so that we have $\partial\tau_{1}(x)=R^{q_{1}}F(x)$. For each $1$-cell $F$ of $\tilde{X}_{1}(q_{1})$ contained in $E$, let $l=d_{\infty}(F,\{v_{E}\})$ and let $\{B_{i}^{F}\}_{i\in I_{F}}$ be an $(2N+1,\delta_{1})$-admissible family which covers $B(x_{l+1},2\delta)\cup\bigcup_{i\in I_{E_{1}}}U_{i}^{E_{1}}\cup\bigcup_{i\in I_{E_{1}}}U_{i}^{E_{2}}$, which exists by Lemma 3.5 for $\delta_{1}=12\delta$. For $F^{\prime}=T_{EE^{\prime}}(F)$, define $\{B_{i}^{F^{\prime}}\}_{i\in I_{F^{\prime}}}=\{B_{i}^{F}\}_{i\in I_{F}}$. Then we have $\operatorname{\mathbf{M}}(\tau_{1}(x))\leq 15L\frac{\varepsilon}{\delta}$ and if $x,y\in F^{\prime}$ then
@@
is supported in $\bigcup_{i\in I_{F^{\prime}}}B_{i}^{F^{\prime}}$. In case $|\mathcal{E}|=1$, we just define the admissible family to be $\{B(x_{l+1},2\delta)\}$. Thus properties (A1) to (A4) hold with $c(k,1)=15$, $n(k,1)=1$, and $K(N,k,1)=12$.
Inductive step 2. Assume $\tau_{j-1}$ has been defined in $\tilde{X}_{j-1}(q_{j-1})_{0}$ ($q_{j-1}=q(k,j-1,p,M,\delta)$) verifying properties (A1) to (A4). Define
$d_{j}$ $=q(k+1,j-1,M,K(N,k,j-1)\delta),$
@@
and for each $j$-cell $F\subseteq E$ of $X_{j}(q_{j})$ the family $\tau_{j}:V(F)\to\mathcal{Z}_{k}(M)$ is localized in a $(K_{2}(N,k,j),K_{2}(N,k,j)\delta)$ admissible family by Lemma 3.20 with $r=\delta$ which yields $K_{2}(N,k,j)=12(K_{1}(N,k,j)+1)$. If the equivalence class $\mathcal{E}$ of $E$ has more than one element, there exists a unique cell $C_{E}$ of $X$ such that $v_{E}\in\operatorname{Center}(C_{E})$ and it holds that $T_{EE^{\prime}}(v_{E})\in\operatorname{Center}(C_{E})$ for every $E^{\prime}\in\mathcal{E}$. Let us enlarge the previously constructed admissible family to a $(K_{2}(N,k,j)+N,3(K_{2}(N,k,j)+1)\delta)$-localized family which also covers $\{U_{i}^{C_{E}}:i\in I_{C_{E}}\}$ (using Lemma 3.5). This will be the family $\{B_{i}\}_{i\in I_{F}}$ and thus we set $K(N,k,j)=3(K_{2}(N,k,j)+1)$. Notice that $\tau_{0}(v_{E})-\tau_{0}(T_{EE^{\prime}}(v_{E}))$ is supported in $\bigcup_{i\in I_{C_{E}}}U_{i}^{C_{E}}$ and therefore in $\{B_{i}\}_{i\in I_{F}}$ for every $E^{\prime}\sim E$.
Define $\tau_{j}(x)=R^{d_{j}}\tau_{j-1}(\Xi(x))$ for each $x\in X_{j-1}\cap X(q_{j})_{0}$, which extends the previous definition of $\tau_{j}$ along $\partial E$. For $x\in X_{j-1}\cap X(q_{j})_{0}\cap E$ set $\Delta_{EE^{\prime}}^{j}(x)=\tau_{j}(x)-\tau_{j}(T_{EE^{\prime}}(x))$. Let $^{\prime}\in\mathcal{E}, EE$$^{\prime}\neq E$. We extend $\Delta^{j}_{EE^{\prime}}(x)$ to $\operatorname{int}(E)\cap\tilde{X}(q_{j})_{0}$ by
@@
$\tau_{j}(x)=\tau_{j}(T_{E^{\prime}E}(x))-\Delta^{j}_{EE^{\prime}}(T_{E^{\prime}E}(x))$
@@
$\operatorname{\mathbf{M}}(\tau_{j}(x))\leq\operatorname{\mathbf{M}}(\tau_{j}(T_{E^{\prime}E}(x)))+\operatorname{\mathbf{M}}(\Delta^{j}_{EE^{\prime}}(T_{E^{\prime}E}(x)))\leq c(k,j)(\frac{L}{\delta})^{n(k,j)}\varepsilon$
if we set c(k,j)=9c(k+1,j1)c(k,j1)c(k,j)=9c(k+1,j-1)c(k,j-1) and n(k,j)=n(k+1,j1)+n(k,j1)+1n(k,j)=n(k+1,j-1)+n(k,j-1)+1. Additionally, if $x,y\in F^{\prime}\cap\tilde{X}(q_{j})_{0}$ then
$\tau_{j}(x)-\tau_{j}(y)=\tau_{j}(T_{E^{\prime}E}(x))-\tau_{j}(T_{E^{\prime}E}(y))-\Delta^{j}_{EE^{\prime}}(T_{E^{\prime}E}(x))+\Delta^{j}_{EE^{\prime}}(T_{E^{\prime}E}(y))$
which by construction of $\tau_{j}$ along $F$ is supported in $\bigcup_{i\in I_{F}}B_{i}^{F}$. In case $\mathcal{E}$ has only one element, the $(K_{2}(N,k,j),K_{2}(N,k,j)\delta)$-admissible families constructed first already have the desired properties. ∎
@@ 3.4. Approximating discrete families of cycles
There exist constants $\varepsilon_{0}(k,p,M)>0$, $\delta_{0}(k,p,M)>0$, $q(k,p,M,\delta)\in\mathbb{N}$ and $N(k,p)\in\mathbb{N}$ such that the following is true. Let $F:X_{0}\to\mathcal{Z}_{k}(M)$ be an $\varepsilon$-fine family, $\varepsilon\leq\varepsilon_{0}$, $0\leq k\leq n-1$, $\dim(X)=p$. Let $0<\delta<\delta_{0}$. Then there exists a discrete family $F^{\prime}:X(q)_{0}\to\mathcal{Z}_{k}(M)$ such that
@@
    $F^{\prime}$ is $c(k,p)(\frac{L}{\delta})^{n(k,p)}\varepsilon$-fine in $X$ and $(N(k,p),\delta)$-localized in $X(q)$.
We start by fixing for each pair of adjacent vertices $v,w\in X_{0}$ a $(k+1)$-chain $\tau_{vw}=\tau(v,w)$ such that $\operatorname{\mathbf{M}}(\tau_{vw})\leq\varepsilon$ and $\partial\tau_{vw}=F(v)-F(w)$. We let $r$ be a sufficiently small number compared to $\delta$ so that the following constructions yield a $\delta$-localized family (it will be enough to choose $r\leq\frac{\delta}{N(k,p)}$ for the constant $N(k,p)$ which is inductively constructed below). We fix points $p_{1},...,p_{L}$ in the manifold $M$ such that $\{B(p_{l},r):1\leq l\leq L\}$ covers $M$. They can be chosen so that $Lr^{n}\leq c(n)\operatorname{Vol}(M,g)$ for some universal constant $c(n)$ which depends only on the dimension of the manifold $M$. For each cell $C$ of $X$ we denote $V(C)=C\cap X_{0}$ the set of its vertices and we choose a preferred vertex $v_{C}\in V(C)$.
Following the strategy in [GL22], we will inductively define for each $1\leq j\leq p=\dim(X)$ a $\delta_{j}$-localized and $\varepsilon_{j}$-fine map $F_{j}:X_{j}(q_{j})_{0}\to\mathcal{Z}_{k}(M)$, where $q_{1}<...<q_{p}=q$. All of the previous maps will agree with $F$ at $X_{0}$. Given a $j$-dimensional cell $C$ of $X(q)$, the map $F_{j}|_{C\cap X(q_{j})_{0}}$ will be given by an interpolation of the values of $F(v)$ over the vertices $v$ of $C$ which does not increase the mass too much with respect to $\max\{\operatorname{\mathbf{M}}(F(v)):v\in C\cap X_{0}\}$. Let us explain how this interpolation works for $j=1$.
Let $C\in\operatorname{Faces}_{1}(X)$ and let $v,w$ be the vertices of $C$. We are going to choose a certain grid $\mathcal{D}_{C}=\{D_{1},...,D_{L}\}$ on $M$ consisting of $L$ domains with disjoint interiors and piecewise smooth boundaries. Roughly speaking, the idea to interpolate between $F(v)$ and $F(w)$ is to start from $F(v)$ and replace $F(v)\llcorner D_{l}$ by the chain with smallest mass among $F(w)\llcorner D_{l}$ and $F(v)\llcorner D_{l}$ for each l=1,,Ll=1,...,L. This operation is done one domain at a time. Thus we start from $z_{0}=F(v)$ and inductively define $z_{l}=z_{l-1}$ if $\operatorname{\mathbf{M}}(F(v)\llcorner D_{l})\leq\operatorname{\mathbf{M}}(F(w)\llcorner D_{l})$ and
$z_{l}=z_{l-1}-\partial[\tau_{vw}\llcorner D_{l}]$
@@
otherwise. It can be shown that $\tilde{z}_{L}=z_{L}$ and hence this allows to define $F_{1}$ along $C(2l)_{0}$ by $z_{0},z_{1},...,z_{L},\tilde{z}_{L-1},...,\tilde{z}_{1},\tilde{z}_{0}$. This has the desired $\delta$-localization property as $z_{l}-z_{l-1}$ is supported in $D_{l-1}\subseteq B(x_{l-1},2r)$ for every $l$ (respectively $\tilde{z}_{l}-\tilde{z}_{l-1}$). We can formalize the previous in the following way. For each $x\in C\cap X(3L)_{0}$ and each $D\in\mathcal{D}_{C}$, we have a certain vertex $w^{D}_{C}(x)\in\{v,w\}$ such that
(2) $F_{1}(x)=F(v)-\sum_{D\in\mathcal{D}_{C}}\partial[\tau(v,w_{C}^{D}(x))\llcorner D]$
@@
for some family $I_{j}:X_{j}(q_{j})_{0}\to\mathcal{I}_{k}(M)$ with small mass in case we try to generalize (3). It will be convenient to work with (5) because when proceeding inductively from the (j1)(j-1)-skeleton to the $j$-skeleton, the preferred vertex $v_{E}$ of a (j1)(j-1)-face $E$ of $C$ will not always coincide with $v_{C}$ (for example when $v_{C}\notin E)$ and therefore we would need to do a certain adjustment to the $\tau(v_{E},w_{j}^{D}(x))$ in order to have (4) which is not necessary if we adopt (5). This problem already arises when $j=2$, as for $x\in E\subseteq\partial C$ with $v_{E}\neq v_{C}$ we would have